Given: f(x) 2x -1, g(x) x + 2 Find f(g(2) 31 28 11 QUESTION 10 O Type here to search
Answer #1

f(g(2)) = f(2 + 2) = f(4) = 2 * 4^2 - 1 = 31

Option - B is correct.